Determine whether the following functions are continuous at a given point. Show your complete solution.

1. ƒ(𝑥)=3𝑥2−4𝑥+2at𝑥=2

2. ƒ(𝑥)=𝑥2−6𝑥−3at𝑥=4


1
Expert's answer
2022-03-21T02:02:32-0400

For a function f to be continuous, at a given point, then:

  • The function must be defined at a point a to be continuous at that point x = a.
  • The limit of the function f(x) should exist at the point x = a,
  • The value of the function f(x) at that point, i.e. f(a) must equal the value of the limit of f(x) at x = a.


1. "ƒ(𝑥)=3𝑥^2−4𝑥+2 \text{ at }𝑥=2"


(i) Substituting x=2 in f(x):

"f(2) = 3(2)^2 -4(2)+2 =3(4)-4(2)+2 = 6"


(ii) Limit at x=2


"\begin{aligned}
\lim_{x \rightarrow2} ƒ(𝑥) &= \lim_{x \rightarrow2} (3𝑥^2−4𝑥+2)\\
& = 3(2)^2 -4(2)+2\\
&=3(4)-4(2)+2\\
&= 6\\
&=f(2)
\end{aligned}"

"\therefore ƒ(𝑥)=3𝑥^2−4𝑥+2 \text{ is continuous at } 𝑥=2"



2. "ƒ(𝑥)=𝑥^2−6𝑥−3 \text{ at } 𝑥=4"


(i) substituting x=4 in f(x)

"\u0192(4)=(4)^2\u22126(4)\u22123 = 16-24-3 = -11"


(ii) Limit at x =4


"\begin{aligned}
\lim_{x \rightarrow 4} f(x) &= \lim_{x \rightarrow 4} (𝑥^2−6𝑥−3)\\
&= (4)^2−6(4)−3\\
& = 16-24-3\\
& = -11\\
& = f(4)
\end{aligned}"

"\therefore ƒ(𝑥)=𝑥^2−6𝑥−3 \text{ is continuous at } 𝑥=4"